LAB Inc reported a net income of $ 150 million on a beginning book value of equity

of $ 1.2 billion in 2017. The firm paid out $ 60 million in dividends, and bought back $ 15

million of stock during the year.

a. Assuming that the firm’s return on equity and reinvestment rate remain the same in 2017,

estimate the expected growth rate in 2018.

b. The average return on equity for the industry is 15%. If LAB Inc’s return on equity changes

to match the industry average in 2018, estimate the expected growth rate in earnings in

2018.
Answer #1

(a) Net Income = $ 150 million, Dividends = $ 60 million, Stock Repurchase = $ 15 million

Retained Earnings = Net Income - Dividends - Stock Repurchases = 150 - 60 - 15 = $ 75 million

Retention Ratio = RR = (Retained Earnings / Net Income) = (75/150) = 0.5

Book Value of Equity = $ 1.2 billion or $ 1200 million

ROE (Return on Equity) = (Net Income / Book Value of Equity) x 100 = (150/1200) x 100 = 12.5 %

Growth Rate = RR x ROE = 0.5 x 12.5 = 6.25 %

(b) If ROE changes to match the industry average of 15 %, the new growth rate would be (15 x 0.5) = 7.5 %